OpenStudy (anonymous):

Solve the equation for all t ∈ [0, 2π) sin 2x + 2 cos x + √3 sin x + √3 = 0 Give the exact sum of the solutions found above.

OpenStudy (anonymous):

sin 2x + 2 cos x + √3 sin x + √3 = 0 2sin x cos x + 2 cos x + √3 sin x + √3 = 0 2cosx (sinx +1) + √3 ( sinx +1) =0 ( 2cosx + √3) (sin x +1) =0 ( 2cosx + √3) =0 or (sin x +1) =0 cosx = -√3 /2 or sin x =-1 Can you solve them from here?

OpenStudy (anonymous):

@shelovespiano ?

OpenStudy (anonymous):

I'm trying to figure this out, i guess... I haven't been taught this and we were expected to figure it out from the book's description. So, from there you'd take the arccos and arcsin then? Or just check a chart?

OpenStudy (anonymous):

Either of them will do, since they are the special angles

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

Could you explain how you got from step three to step 4 please?

OpenStudy (anonymous):

cosx = -√3 /2 x = 180 - 30 =? or x = 180 + 30 =? sin x =-1 x = 270 Put the angle back to the equation to check if it satisfies the equation

OpenStudy (anonymous):

Do factorization, there is a common factor (sinx+1) in the 2 terms

OpenStudy (anonymous):

OH, I see now. thanks again!

OpenStudy (anonymous):

Welcome :)